OpenStudy (lilkg77):

The first quartile of a data set is 41, and the third quartile is 57. Which of these values in the data set is an outlier? A. 83 B. 73 C. 19 D. 39

OpenStudy (lilkg77):

@wolfcherry

OpenStudy (lilkg77):

@mathmale

OpenStudy (welshfella):

the interquartile range is 57-41 = 16 Any point outside 1.5 times the interquartile the lower or upper quartile is an outlier so 1.5 * 16 = 24

OpenStudy (welshfella):

So which points are outsdie tihs rangge

OpenStudy (anonymous):

RIGHT

OpenStudy (welshfella):

so 41 - 24 = 17 - so 19 is NOT an outlier can you see which ones are?

OpenStudy (welshfella):

Hint: there is one above the upper quartile (which is 57)

OpenStudy (welshfella):

add 24 to 57 and see which ones are outside the result

OpenStudy (lilkg77):

81

OpenStudy (welshfella):

ok so what have we got above 81?

OpenStudy (welshfella):

- what number in the list is > 81

OpenStudy (lilkg77):

83

OpenStudy (welshfella):

yes So that is an outlier

OpenStudy (welshfella):

any number less than 17 or greater than 81 is an outlier

OpenStudy (texaschic101):

Q3 - Q1 = IQR Q1 - 1.5 * IQR....anything below this is an outlier Q3 + 1.5 * IQR....anything above this is an outlier

OpenStudy (lilkg77):

Thanks for the help @welshfella @texaschic101

OpenStudy (welshfella):

yw

OpenStudy (texaschic101):

yw